OpenStudy (anonymous):

what are three consecutive numbers whose sum is 138 by writing and solving an equation

OpenStudy (oaktree):

Write (x)+(x+1)+(x+2)=138 and solve.

OpenStudy (anonymous):

this the answer

OpenStudy (oaktree):

So 3x+135 Therefore x=135/3 or 45

OpenStudy (oaktree):

So the three numbers are x,x+1, and x+2 which gives 45, 46, and 47.